OpenStudy (anonymous):

What effect does doubling the height of a cone have on the volume of the cone if the radius remains the same? a. The volume is doubled. b. The volume is quadrupled. c. The volume is halved. d. The volume is multiplied by one-forth.

OpenStudy (tkhunny):

Compare: \(\dfrac{1}{3}\pi r^{2}h\) and \(\dfrac{1}{3}\pi r^{2}(2h)\)